OpenStudy (anonymous):

Write an equation of the line that is perpendicular to 2x + 4y = 6 and passes through the point (6,4).

OpenStudy (sesca89):

y = -.5x - 1.5 graph it and check...

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

Write an equation of the line that is perpendicular to 2x + 4y = 6 and passes through the point (6,4). A) y = 2x + 4 B) y = 2x - 8 C) y = -2x + 16 D) y = - 1

OpenStudy (anonymous):

sorry D. is messed up but it is y=-1

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

first you wanna do is put it in slope-intercept form which should be 4y=-2x=6 perpendicular have the negative repicals so -2x should be 1/2 for your slope so you should be @ 4y=1/2x=b becuase you dont know what b is plug in your (x,y) so 4(4)=1/2(6)=b and that should be 16=3+b move the 3 over 16-3+b b=13 then plug in Y=1/2x+13

OpenStudy (anonymous):

Oh My Goodness Thank You Soo Much!